OpenStudy (anonymous):

Solve the system of equations and choose the correct answer from the list of options. d + e = 12 −d + e = −8 Label the ordered pair as (d, e). (0, 0) (2, 10) (10, 2) (12, 0)

OpenStudy (anonymous):

c

OpenStudy (anonymous):

please help ill fan and medal

OpenStudy (anonymous):

10,2

OpenStudy (anonymous):

thxx

OpenStudy (kidrah69):

a negative + positive cancel out so you are left with e=12/e=-8

OpenStudy (anonymous):

help with 1 more?

OpenStudy (anonymous):

10+2=12

OpenStudy (anonymous):

)A system of equations is given below: y = 4x − 3 2x + 7y = 41 Which of the following steps could be used to solve by substitution? 2(y = 4x − 3) + 7y = 41 4x − 3 = 2x + 7y 2x + 7(4x − 3) = 41 2(y = 4x − 3)